Keith received an excellent interest rate for his loan of $12,000. Calculate the nominal interest rate compounded semi-annually if the loan accumulated to $12,642.11 in 5 years and 6 months


1
Expert's answer
2021-02-24T07:00:08-0500

"A=P(1+r\/m)^{n}"

"m=2, n=2\times 5+1=11"

"12642.11=12000(1+r\/m)^{11}"

"\frac{12642.11}{12000}=
(1+r/2)^{11}"

"1.0535=(1+r\/2)^{11}"

"1.0535^{\frac{1}{11}}=1+r/2"

"1.0047=1+r\/2"

"0.0047=r\/2"

"r=0.0094"